Xy^2z^3=8 at (2,2,1)


1
Expert's answer
2021-10-25T16:33:57-0400

Let "f(x,y,z): xy^2z^3-8=0"

"\therefore \frac{\delta f}{dx}=y^2z^3; \frac{\delta f}{dy}=2xyz^3; \frac{\delta f}{dz}=3xy^2z^2"

At "(2,2,1):"

"\frac{\delta f}{dx}=4;\frac{\delta f}{dy}=8;\frac{\delta f}{dx}=24"

Therefore, equation of tangent is:

"4(x-2)+8(y-2)+24(z-1)=0\\
\Rightarrow x-2+2y-4+6z-6\\
\Rightarrow x+2y+6z=12"